# A Comprehensive Survey of Mixture-of-Experts: Algorithms, Theory, and Applications

### v1

@@ V-A Computer vision
GANs [266]
are known to struggle to learn multimodal data distributions when processing complex datasets, leading to generated images with poor-quality. To address this, [156] proposes MEGAN with multiple generator networks, where each network focuses on learning specific modal distributions in the dataset. This design allows each generator to focus on different data subsets, generating more diverse and higher-quality images.
@@ V-B Natural Language Processing
When speech is transcribed into text, errors from Automatic Speech Recognition (ASR) systems [278]
often affect the accuracy of subsequent NLU components. To address this, [159] proposes an MoE-based framework, namely MoE-SLU, aiming to mitigate the impact of ASR errors on spoken language understanding performance. MoE-SLU employs three strategies to generate additional transcripts and uses MoE to weight and average these transcripts. By enhancing the model’s ability to capture keywords, MoE-SLU is more robust to ASR errors, achieving state-of-the-art performance on three benchmark SLU datasets. The performance of MoE-SLU can be further improved through regularized predictions.
@@
2) Natural Language Generation. NLG can transform structured data or concepts into natural and fluent text, widely applied in machine translation [279], dialogue systems [280], content creation [281], and information summarization [282], among other fields
. To achieve diverse and precise language expression, NLG systems require strong model capacity to capture complex linguistic structures and contextual dependencies, while effectively handling interference in multi-task learning to maintain high performance across different scenarios. MoE allows models to dynamically select the most suitable expert networks and flexibly adjust model capacity, reducing task interference and enhancing model specialization, thereby contributing to the construction of more powerful and flexible NLG systems.
@@ VI-E Tailored algorithm design
MoE models have shown great promise in various machine learning paradigms, but their potential in combination with other learning paradigms remains underexplored. Future work should investigate the integration of MoE with other learning frameworks, such as contrast learning, transfer learning, and self-supervised learning. For example, in federated learning, MoE models could be used to handle heterogeneous data distributions across clients by assigning different experts to different clients. Exploring these combinations could lead to more versatile and powerful learning systems.

### v2

@@ V-A Computer vision
GANs [285] are known to struggle to learn multimodal data distributions when processing complex datasets, leading to generated images with poor-quality. To address this, [169] proposes MEGAN with multiple generator networks, where each network focuses on learning specific modal distributions in the dataset. This design allows each generator to focus on different data subsets, generating more diverse and higher-quality images.
@@ V-B Natural Language Processing
When speech is transcribed into text, errors from Automatic Speech Recognition (ASR) systems [297] often affect the accuracy of subsequent NLU components. To address this, [172] proposes an MoE-based framework, namely MoE-SLU, aiming to mitigate the impact of ASR errors on spoken language understanding performance. MoE-SLU employs three strategies to generate additional transcripts and uses MoE to weight and average these transcripts. By enhancing the model’s ability to capture keywords, MoE-SLU is more robust to ASR errors, achieving state-of-the-art performance on three benchmark SLU datasets. The performance of MoE-SLU can be further improved through regularized predictions.
@@
2) Natural Language Generation. NLG can transform structured data or concepts into natural and fluent text, widely applied in machine translation [298], dialogue systems [299], content creation [300], and information summarization [301], among other fields. To achieve diverse and precise language expression, NLG systems require strong model capacity to capture complex linguistic structures and contextual dependencies, while effectively handling interference in multi-task learning to maintain high performance across different scenarios. MoE allows models to dynamically select the most suitable expert networks and flexibly adjust model capacity, reducing task interference and enhancing model specialization, thereby contributing to the construction of more powerful and flexible NLG systems.
@@ VI-E Tailored algorithm design
MoE models have shown great promise in various machine learning paradigms, but their potential in combination with other learning paradigms remains underexplored. Future work should investigate the integration of MoE with other learning frameworks, such as contrastive learning, transfer learning, and self-supervised learning. For example, in federated learning, MoE models could be used to handle heterogeneous data distributions across clients by assigning different experts to different clients. Exploring these combinations could lead to more versatile and powerful learning systems.